solve ( 3x y^3 + 2y )dx + (2x^2 y + x ) dy = 0


1
Expert's answer
2021-04-29T16:47:01-0400

Given equation is-

"( 3x y^3 + 2y )dx + (2x^2 y + x ) dy = 0"


This equation is of the form Of exact differential equation-

"Mdx+Ndy=0"


So The solution of the equation is-


"\dfrac{3y^3x^2}{2}+2yx+0=0"


"3y^3x^3+4yx=0"